### v3

# DropleX: Liquid sensing on tablet touchscreens

Abstract.
We present DropleX, the first system that enables liquid sensing using the capacitive touchscreen of commodity tablets. DropleX detects microliter–scale liquid samples, and performs non-invasive, through-container measurements to detect whether a drink has been spiked or if a sealed liquid has been contaminated. These capabilities are made possible by a physics-informed mechanism that disables the touchscreen’s built-in adaptive filters, originally designed to reject the effects of liquid drops such as rain, without any hardware modifications. We model the touchscreen’s sensing capabilities, limits, and non-idealities to inform the design of a signal processing and learning-based pipeline for liquid sensing. Our system achieves 96–99% accuracy in detecting microliter-scale adulteration in soda, wine, and milk, 93–96% accuracy in threshold detection of trace chemical concentrations, and 86–96% accuracy in through-container adulterant detection. Given the predominance of touchscreens, these exploratory results can open new opportunities for liquid sensing on everyday devices.
1. Introduction
Imagine testing whether your drink has been spiked at a bar by simply setting the cup on your tablet screen. Or imagine placing a single drop of water, soda, or milk on your touchscreen to know if it has been contaminated. Finally, imagine detecting trace concentrations of biomolecules such as DNA. While touchscreens have conventionally been designed to reject the effects of liquids such as raindrops, treating it as unwanted noise, this paper takes the opposite approach and investigates if instead commodity capacitive touchscreens can be repurposed to perform liquid sensing. Given the predominance of capacitive touchscreens on modern computing devices, the ability to perform liquid sensing on touchscreens could transform tablets into a portable “lab-on-a-pad” as an accessible means of adulteration and concentration detection in everyday settings, resource-constrained environments and classrooms.
Conventional methods (con, 2025c, a, b) for performing capacitance-based liquid sensing typically rely on a dedicated device with large probes designed to be immersed into open cup solutions to a minimum depth, and are typically limited to measuring a single sample at a time. While there is rich prior work (Yeo et al., 2016; Rahman et al., 2016; Wang et al., 2017; Dhekne et al., 2018; Yue and Katabi, 2019; Ha et al., 2020; Yun et al., 2024; Chan et al., 2022b, a) in the mobile and wireless community on liquid based sensing using radar, RFID, vibration motors and cameras, including work leveraging an external standalone capacitive touchscreen, they are largely focused on sensing larger quantities of liquid, require some form of dedicated, specialized hardware and non-trivial hardware modification to commodity devices, or do not demonstrate the ability to classify between different liquids.
We introduce DropleX, the first system to enable liquid sensing on the integrated capacitive touchscreen of commodity tablets. Our system is capable of detecting microliter-scale liquid samples and performing non-invasive through-container measurements even for sealed liquids. It is able to conduct multiple tests concurrently across the screen, allowing simultaneous characterization of different samples.
However, repurposing the tablet touchscreen to perform liquid sensing is not straightforward and requires tackling several key challenges. First, commercial touchscreens have built-in adaptive filters that continuously recalibrate capacitance readings to a baseline to reject transient effects such as those caused by liquid drops. As a result, any liquid or container placed on the screen is only registered for a few seconds before the readings fades. To address this, we develop a physics-informed mechanism that can disable the filter temporarily for the measurement without any hardware or software modifications.
Specifically, we observe that the adaptive filter is turned off in the presence of a finger touch event. This is because recalibrating during contact would cause the system to disregard the touch input, and make touch interaction impossible. Our idea is to mimic a permanent touch event, by having the user deposit and draw up a small “priming” drop of water onto the screen which leaves behind a thin conductive film. Unlike a bulk drop of water that confines the electric field to the screen, the thin film acts primarily as a conductor and causes part of the electric field to be shunted to parasitic ground and is effectively recognized as a finger. This causes the adaptive filter to be disabled and allows for reliable measurements of liquids placed on the screen.
This deposition and removal process can be performed using common applicators including eye droppers, and can also be performed without any special equipment. The user can simply dip their finger into a cup of water, allowing a small pendant drop to adhere to the skin, and then touch the screen to deposit it. The drop can then be drawn up by lightly touching it with a tissue, which wicks it away through capillary action.
Second, the measured capacitance values for different liquids or concentrations can overlap, making it difficult to distinguish using a single scalar metric. To overcome this, we instead leverage the spatial spreading patterns of the liquid, as well as the fringing fields from the electric field lines that extend beyond the drop boundary, and leverage a learning-based approach to capture pertinent features that can accurately characterize the drop.
Third, as touchscreens are not designed for precise measurements of capacitance for liquid sensing, the electrode grid exhibits spatially varying sensitivity and time-varying noise across the screen. To address this challenge, we develop a one-time, multi-point calibration method to generate a sensitivity compensation map over the screen, and apply temporal averaging to reduce noise.
Our contributions are as follows:
We present DropleX, the first system that enables liquid sensing on commodity tablet touchscreens without any hardware modifications. We demonstrate adulteration detection accuracies of 96–99% for soda, wine, and milk samples, detection of trace amounts of biochemicals including DNA, NaCl, IPA with accuracies of 93–96%, and through-container sensing with accuracies of 86–96% across multiple containers.
We run characterization and benchmark experiments across different sample volumes, temperatures, as well as the effect of interfacial layers such as thin-films and different containers on the capacitance readings.
We develop a physics-informed model that relates the device’s measurements to physical capacitance readings.
We develop an end-to-end software infrastructure, including utilities to read the capacitance heatmaps, a real-time Android tablet app and desktop app to visualize the readings. We plan to open-source the code and tutorial for enabling this capability on tablets as well as our dataset upon publication as a key technical primitive that the community can build on to enable future sensing systems on touchscreens.
As the proposed approach involves measurements of capacitance heatmaps already accessible to technology manufacturers in software, it can in principle be deployed through a software or firmware update to commodity devices.
2. Related work
| Reference | Technology | Quantity measured | Ubiquitous smart device | No hardware additions or modifications needed | Distinguishes between concentrations |
| Macroscopic-scale methods () | |||||
| RadarCat (2016) (Yeo et al., 2016) | Radar | Reflected signal | ✗ | N/A | ✗ |
| Nutrilyzer (2016) (Rahman et al., 2016) | Photoacoustic effect | Photoacoustic spectra | ✓ | ✗ | ✓ |
| TagScan (2017) (Wang et al., 2017) | RFID | RSS, phase change | ✗ | N/A | ✓ |
| LiquiID (2018) (Dhekne et al., 2018) | Ultra-wideband radios | Permittivity | ✗ | N/A | N/A |
| CapCam (2019) (Yue and Katabi, 2019) | Vibration motor, camera, ripples | Surface tension | ✓ | ✓ | ✓ |
| RF-EATS (2020) (Ha et al., 2020) | RFID | Amplitude and phase change | ✗ | N/A | ✓ |
| Microliter-scale methods () | |||||
| Prior touchscreen systems (2016–2021) (Horstmann et al., 2021; Won and PARK, 2012; Won et al., 2016; Lee et al., 2018) | Capacitive sensing | Conductivity | ✗ | ✗ | ✗ |
| Chan et al. (2022) (Chan et al., 2022b) | LiDAR, laser speckle, camera | Viscosity | ✓ | ✗ | ✓ |
| Chan et al. (2022) (Chan et al., 2022a) | Vibration motor and camera | Viscosity | ✓ | ✓ | N/A |
| DropleX (ours) | Capacitive sensing | Capacitance | ✓ | ✓ | ✓ |
Prior work can be broadly divided into three different domains (Table 1).
Touchscreens for liquid sensing. Several works (Horstmann et al., 2021; Won and PARK, 2012; Won et al., 2016; Lee et al., 2018) have proposed the use of capacitive touchscreens to measure conductivity properties of different electrolytes, polymers, and biomolecules including DNA such as of the H1N1 virus, Chlamydia trachomatis. However, they suffer from several key limitations: (1) they use an external standalone display which is not part of a smart device, and do not face or address the practical challenge of an adaptive filter (2) they uses an electrode array pitch (separation) that is 2 to 3.5 times smaller than what is found in conventional tablet touchscreens, (3) it coats the touchscreen surface with a protective polymer of parylene C, which is impractical for everyday use. Collectively, these constraints fail to capture the practical realities of performing liquid sensing on commodity touchscreens integrated into tablets, and therefore do not encounter or address the range of real-world challenges our work considers. Moreover, prior efforts do not attempt to classify between different liquid types or concentrations, and their reported capacitance values often exhibit significant overlap across samples.
Our work differs in three important aspects: (1) It performs liquid sensing directly on unmodified tablet touchscreens, and overcomes the challenges of hardware filters and other non-idealities to make this possible. (2) It goes beyond reporting raw measured capacitance values and develops a pipeline to perform liquid classification for different classes of liquids. (3) Importantly, we demonstrate through-container sensing scenarios for a variety of containers, which enables use-cases in everyday settings outside lab testing of liquids.
Touchscreens for novel interaction techniques. Beyond liquid sensing, there has also been interest in leveraging smart device touchscreens for enabling novel and tangible interaction capabilities such as by detecting ungrounded conductive objects on the screen (Voelker et al., 2013, 2015; Rekimoto, 2002), applying super-resolution to detect the visual signature of the object through motion (Mayer et al., 2021), estimating finger contact angle (Xiao et al., 2015) detecting between fingertips, raindrops and water smear (Tung et al., 2018), and differentiating between different user’s handprints for authentication (Guo et al., 2015).
A key challenge with enabling this is that commercial touchscreens specifically implement an adaptive filter to detect fingers which are electrically grounded, and reject other objects. Various works (Jansen et al., 2012; Kratz et al., 2011; Yu et al., 2011a, b) have presented alternative approaches to disabling this filter such as having the body capacitance of the user couple to the touchscreen such as through a conductive element connected to the screen (Wiethoff et al., 2012), or using a conductive element connected to a relatively grounded object such as the battery ground connector of the tablet or a second area of the display as ground (Voelker et al., 2013). The challenge with this approach is that it uses dedicated custom conductive markers, and the grounding is dependent on the geometry of the device. In contrast, our approach leverages the existing benign liquids such as water which is ubiquitous as a means of tricking the phone into thinking there is a grounded object in contact with the screen and overriding the adaptive filter. We note further that none of these systems are designed for liquid sensing and are focused around detecting tangible objects or widgets on the screen. Our work is focused on sensing the properties of micro-liter-scale samples of liquid based on mutual capacitance readings of the touchscreen.
Liquid-based sensing. There is rich prior work in the space of liquid sensing using mobile and wireless systems ranging from using radar-based approaches including UWB (Dhekne et al., 2018), RFID (Wang et al., 2017; Xie et al., 2019; Ha et al., 2020, 2018), mmWave radar (Yeo et al., 2016), photoacoustic-based (Rahman et al., 2016) approaches for liquid contaminant detection. These systems offer complementary designs to the touchscreen-based approach and are designed to operate on liquid samples that are at least 10 mL which is three orders of magnitude larger than what touchscreen-based approaches can measure. These systems are also focused on using custom, dedicated hardware systems which can be challenging to scale outside laboratory settings in particular to resource-constrained environments.
Prior work using mobile devices for liquid sensing includes CapCam (Yue and Katabi, 2019) which leverages the vibration motor on the smartphone placed on top of a cup to vibrate the liquid and measure the wavelength of the capillary waves and compute surface tension. The constraint of this approach is that the sample needs to be placed in a cup that is filled to a depth of 25 to 45 mm so that the focal length of the camera lens matches the depth of the liquid, which excludes it from practical use of analyzing smaller liquid samples. Chan et al. (Chan et al., 2022a) measured blood clotting time on smartphones for 10 samples of blood and plasma using a vibration motor, camera, and a tiny copper particle that needs to be deposited into each sample. A similar related work (Chan et al., 2022b) also performed blood clot testing using the LiDAR sensor and laser speckle to measure the viscosity of 10 samples of blood, milk, and other household liquids. However, these systems require custom hardware components or modifications: specialized copper particles, or smartphone cameras with the near-infrared filter removed to analyze the LiDAR-based laser speckle patterns. They are also focused on sensing different liquid properties: surface tension and viscosity, which rely on mechanical sensing as opposed to electrical sensing, which is what our work focuses on.
Our work is complementary to these approaches in spirit in that it is focused on leveraging commodity smart devices for liquid sensing, but it does not require any hardware modification or special hardware components to perform liquid sensing.
3. Primer on capacitive sensing on touchscreens
Modern multi-point touchscreens (Fig. 2a) are comprised of multiple layers starting from bottom to top:
Display: Ranges from liquid crystal to light emitted diode based displays.
Insulating substrate: Provides mechanical structural stability, generally glass.
Driving lines: This is an electrode film, often polyethylene terephthalate (PET), with transparent electrode lines coated typically with indium tin oxide (ITO) which is known for its electrical conductivity, optical transparency, resistance to moisture, and ability to be deposited as a thin film. The touchscreen controller sends a driving AC signal along these.
Dielectric layer: Usually PET or an optically clear adhesive (OHA) to mechanically bond and electrically isolate the top and bottom two ITO layers, prevents shorting between them.
Sensing lines: Similar to the driving line, but runs orthogonal or opposing to it to created a grid. It senses how much of the AC signal couples through the dielectric. The two electrode layers and the dielectric as a capacitor.
Protective covering: Usually glass with anti-reflective coating, which insulates the sensors from direct touch.
Touchscreens operate on the idea of mutual capacitance which forms between the two electrode lines (Fig. 2b) (Barrett and Omote, 2010; Hessar et al., 2016). The screen structure is composed of a 2-D array of electrodes from the driving and sensing lines. On the driving line is an AC input drive signal , which creates a shared electric field between the layers, this is sensed by the other layer with the sensing electrodes. We show in Fig. 2c the driving signal of Samsung Galaxy Note 10.1 captured on an oscilloscope (T3DSO1302A, Teledyne LeCroy). The signal is captured by placing the oscilloscope probe directly in contact with the top of the touchscreen and connecting the ground to a metal pole. The driving frequency is approximately 100 kHz with a peak to peak distance is approximately 2.4 mV.
In a touchscreen without any object in contact with it, a mutual capacitive field forms between the driving and sensing electrodes (Fig. 3a). When a conductive object is near, it disturbs the shared electric field. In the case of grounded conductive objects like a human finger that is coupled to Earth ground, the electric field lines are pulled towards the finger and it reduces the capacitance between two electrodes at an intersection (Fig. 3b). The touchscreen controller is able to read the change in capacitance and detect where it has occurred. The human body serves as an effective ground because it is capacitively coupled to Earth, forming a large conductive body with a low-impedance pathway to ground. For the electric field, this provides an easier route to discharge, as the finger is in closer proximity to the screen compared to some of the sensing electrodes which can be farther away and have higher impedance.
In contrast, when placing a liquid sample on the screen it alters the electric field coupling between the driving and sensing electrodes in two ways (Fig. 3c):
Conductivity effect. Liquids containing free ions (e.g. salt water) act similar to the finger in that it can conduct small amounts of current. This provides a partial path to ground as it is parasitically coupled to the air and the device chassis itself. This lossy coupling causes part of the AC signal to leak through the liquid’s conductivity. Under the parallel-plate capacitor approximation, the effective conductive contribution to capacitance can be expressed as:
(1) where is electrical conductivity of the liquid, is the electrode area, is the separation distance or pitch between electrodes, and is the angular frequency of excitation . The remaining charge is redistributed to neighboring electrodes, effectively reducing the mutual capacitance readings (Zimmerman et al., 1995).
Permittivity effect. When liquid with a relative permittivity () higher than air contacts the screen, it acts as a dielectric that increases the degree of polarization in the electric field region. Specifically, what happens is that polar molecules align partially with the field, storing more electric energy and increasing the effective capacitance between electrodes even if the liquid is not strongly conductive according to:
(2) where is vacuum permittivity. Non-conductive or poorly conductive liquids such as oils and alcohols influence the measured capacitance primarily through this effect creating an increase in the measured capacitance.
Overall, as a liquid sample can act as both a conductor and a dielectric and the measured capacitance on the display is both a function of both the liquid’s conductivity and permittivity:
Building on these principles of capacitive sensing, the next section describes how they can be leveraged to enable liquid sensing on touchscreens.
4. Enabling liquid sensing on touchscreens
Effect of touchscreen’s liquid rejection filter. While it may appear straightforward to read the capacitive readings from the tablet to perform liquid sensing an immediate challenge becomes apparent when depositing a drop of water on the screen: the capacitive readings fade away. We demonstrate this effect in Fig. 4a which shows the capacitance measured using our tablet’s reported digital units over a period of 25 seconds for two drops of tap water each with a volume of 500 . The capacitance readings are logged at our device’s maximum polling rate of 1.67 Hz corresponding to a frame measured every 0.6 s, implementation details provided in Sec. 7.
The plot reveals two observations, first there is an initial sharp drop in the capacitance readings (using device units). Second, it shows that the measured capacitance reverts back to baseline. For sample 1, it reverts back to baseline quickly by the next frame, while for sample 2, this change was more gradually, taking approximately 18 s.
The reason for this is that manufacturers specifically design software heuristics known as adaptive filters that actively reject liquids such as rain and other spurious events that change the capacitance readings. The ability to disable this adaptive filter from constantly recalibrating to a baseline could enable our system to measure the effect of liquid samples on the capacitance values more reliably.
We observe that the adaptive filter is disabled when a finger, the intended input for touchscreens, is in contact with the display. Specifically, the measured capacitance remains stable over time. This is expected behavior, since the touchscreen must maintain a persistent signal for a valid touch, and recalibrating the finger’s contact back to baseline would cause it to disappear and make normal touch interactions impossible.
Mimicking a permanent touch to disable the filter. It is of course, however, impractical to expect the user to constantly place their finger in contact with the screen throughout the measurement. Our key idea is to mimic a permanent touch event to disable the adaptive filter and enable liquid sensing. Our approach is based on our prior modeling of how a liquid sample affects the touchscreen’s capacitive field (Fig. 3). Specifically, when a liquid is deposited in the screen both its conductivity and permittivity change the capacitive field. When a drop of water is deposited on the screen, the dielectric permittivity term dominates compared to conductance. This is because a sample of liquid is thick, and in the case of water can be 2 mm in height due to its surface tension, which can be thicker than tablet touchscreens which can have a thickness of 1 mm (thi, 2025). Because of this the electric field is largely confined to within the liquid.
If there was a way to modify the liquid so that the conductivity term dominates, it could then mimic the effect of a finger. Specifically, if the liquid height could be reduced to a thin film for the electric field to pass through, the conductivity term would dominate and it would function like a finger. Our key observation is that when drawing up the liquid, a thin film remains which causes the electric field to be shunted to parasitic ground, and is effectively recognized by the screen as a “finger”.
To demonstrate this, we deposit a 500 “priming drop” of water onto an area of the screen which initially induces a positive change in measured capacitance (negative in device units). After this, however upon drawing up the sample using the pipette the capacitance values become negative (positive in device units) and causes a region of pixels to be in a permanently “stuck” position, effectively emulating a permanent touch. The effect on the measured capacitance is illustrated in Fig. 5b where it can be observed that depositing this initial sample results in a sharp drop in capacitance of around -200 device units, and after it is drawn up, it reaches a stable a value of approximately 400 device units.
After this, the adaptive filter is disabled, and the liquid sample of interest can be deposited elsewhere on the screen and its effects can be examined. We show in Fig. 4b that after disabling the effects of the adaptive filter, we deposit two samples of tap water as before. This results in a sharp drop in mutual capacitance, and a transient stabilization period of about one frame, after which the sample readings remain stable across the measurement period of 25 s. We evaluate the stability of these readings over a longer period in Sec. 8.
Importantly, we note that a pipette or specialized applicator is not required to deposit or draw up the priming drop. As show in Fig. 6, the user can simply dip their finger into water, where surface tension forms a small pendant droplet that remains attached to the skin, and gently touch the touchscreen to transfer it onto the surface. The same droplet can then be immediately removed using a regular tissue, which draws the liquid away via capillary wicking as the water is absorbed by its fibers.
We note that prior works (Voelker et al., 2013) for tangible user interfaces which place ungrounded conductive objects on the screen (e.g. copper tape, coins, keys) have noted this effect and face a similar challenge. These prior works propose including a small conductive clip to ground a point on the screen through capacitive coupling with the chassis of the tablet or as a bridge between electrodes on the screen in order to permanently ground it. However, this requires a custom clip or conducting bridge. Our proposed approach relies on just using tap water which is ubiquitous.
Capacitance Heatmap Signatures of Liquids. We show in Fig. 7 an example capacitance heatmap measured after disabling the adaptive filter for 200 drops of tap water and isopropyl alcohol. The plot shows that different liquids are able to produce different signatures which can be captured by the (1) spatial spreading patterns, (2) capacitance values at the centroid and adjacent cells, (3) fringing patterns around the drop which occur because electric field lines between electrodes extend beyond the droplet boundary and couple to the liquid. We note these fringing patterns are more pronounced in liquids with higher permittivity such as tap water.
5. Modeling capacitance measurements
With the filter disabled, we now quantify how liquid properties of conductivity and permittivity contributed to the measured capacitance values.
Characterization of Liquid Concentration. We perform an experimental analysis by placing different concentrations of salt water (NaCl) and isopropyl alcohol (IPA) concentrations of deionized water on the touchscreen and measured the device’s capacitance readings. Specifically, we used alcohol concentrations of 0, 20, 40, 60, 80, 100% and salt water concentrations of M. The reason why we selected these two liquids is because increases in alcohol concentration decreases the liquid’s permittivity while its conductivity is largely constant, while conversely increases in salt water concentration will increase its conductivity while its permittivity is relatively constant. This allows us to decouple the effects of conductivity and permittivity on the measured capacitance and separately analyze their influence on the touchscreen’s response.
We show in Fig. 8a,b the conductivity of each sample as measured using an off-the-shelf electrical conductivity meter (ec_, 2025), as well as permittivity values obtained from related literature (Park et al., 2006). We note that we do not obtain the permittivity values experimentally, as this typically requires a microwave network analyzer (vna, 2025b) with a dielectric probe (vna, 2025a) costing over a hundred thousand dollars such as used in prior liquid sensing works (Dhekne et al., 2018), and is prohibitively expensive to acquire. We can observe in the table that the capacitance readings from the NaCl and IPA dataset also span non-overlapping ranges.
We show in Fig. 8a,b the measured capacitance reading as the IPA and NaCl concentrations increases. For each liquid we use drops of volume 500 and perform ten replicates. We can observe a general trend where the centroid value in device units increases with IPA concentrations, while it decreases and plateaus with NaCl concentrations. To further explore how liquid composition affects the capacitance response we examined aqueous solutions of DNA sodium salt from calf thymus (dna, 2025) at different concentrations of 0, 20, 40, 60, 80 ng/, and show that the centroid capacitance values in device units decreases (Fig. 8c).
We note here that while the mean values do show a trend for these three liquids, the error bars overlap, suggesting that using a single metric such as the centroid value is not sufficient for classifying between different concentrations, and leveraging additional spatial information is needed.
Physics-Informed Model. We aim to fit a model that explains these trends, specifically focusing on the data points for the concentrations of IPA and NaCl. To do this, we first leverage a physics-informed model based on the idealized parallel-plate capacitor model Eq. 3. However, this idealized model alone is insufficient for two reasons:
The touchscreen reports capacitance in arbitrary digital units, without a clear mapping to physical units of farads.
- (2)
To model these effects we assume that the mapping for touchscreen units is a linear mapping and assume a linear scaling factor and offset . While for the fringing effects we assume it is dominated by a quadratic term . Ultimately we model the capacitance as:
When fitted to the alcohol and NaCl water dataset, we find that Pearson’s correlation coefficient is (Fig. 9). The fitted values are , , , .
We consider an alternative modeling approach which is entirely empirical and leverages a weighted combination of the conductivity and permittivity terms. We experimentally evaluate linear, quadratic, log-transformed, and power law based methods of modeling and find that a quadratic based modeling formulation produces the best fit
The result in Fig. 9 shows that the empirical quadratic based model is able to fit the data with a Pearson’s correlation of with fitted terms , , , , , . This suggests that the empirical relationship between conductivity, permittivity, and measured device capacitance is non-linear, but that a 2nd-order polynomial formulation can characterize this relationship.
6. DropleX design
Now that we are able to perform liquid sensing, in this section we:
6.1. Empirical characterization and analysis
6.1.1. Multi-scale spatial sensitivity analysis
Here, we characterize the spatial sensitivity of the electrodes at the micro-scale, and at the macro-scale across the entire screen.
Micro-scale characterization. Given that electric fields may not be uniform around an electrode, our goal here is to analyze where is the best position around a cell to deposit a drop. To do this, we deposited water samples of 10 and 25 at the center of an electrode, at the corner of the electrode which would be at the intersection of a drive and sense line, along a horizontal line and along a vertical line. We repeat this for 10 droplets for each of the volumes.
To characterize the change in capacitance instead of using the centroid which only measures the value at a single cell, which would make sense if the droplet was deposited only in the center. For drops in other locations which straddle the boundary between multiple cells we sum together the values across cells with a negative change in capacitance. We show in Fig. 11a the magnitude capacitance of the droplet. Both the 10 and 25 samples had comparatively higher capacitance values and lower variation at the center and vertical locations, and lower values at the corner and horizontal locations. The results suggest that placing droplets at the center of the electrode produce the most reliable results. This is the position that we adopt in all experiments. We note however that the use of larger volume samples occupy multiple cells and our system captures features of the spatial information in the region around the centroid of the drop for downstream applications.
Macro-scale characterization. Here, we characterize electrode sensitivity spatially across the tablet screen. These variations are often due to hardware non-idealities such as variations in dielectric thickness, grounding, and shielding around the sides of the screen compared to the center. Tackling this challenge is important to ensure that measurements of the same liquid and volume across different areas of the screen are uniform.
To characterize this, we performed an experiment by depositing a 50 drop of water at the center of each electrode cell at predefined intervals until the screen was covered. We repeated this procedure six times. In these six maps if a measurement occurred at the same pixel location across multiple maps, we averaged the results. In the end, we had a reading for 546 cells across the full grid of 52 32 = 1664 cells.
Given calibration points where is the screen’s response, we generate the interpolated heatmap using a radial basis function with thin plate spline kernel using a smoothing factor .
where is the number of calibration points, is the Euclidean distance between the interpolation point and each calibration point , and the basis function is defined as . The weights are determined by solving the regularized linear system:
with , , and .
The sensitivity map in Fig. 11b, shows two clusters to the left and right of the screen which are sensitive, while the sides and the center have low sensitivity.
The compensation factor in each grid cell is normalized using the function . Here the minimum value across all calibration points corresponds to the most responsive region of the screen. This compensation map represents the inverse of the interpolated sensitivity, and is a small term to prevent numerical instability from division by zero. The compensation is subsequently applied by element-wise multiplying the measured heatmap with the computed compensation map.
6.1.2. Effect of sample volume
We examined the effect of liquid volume on the measured centroid capacitance. We perform this experiment at volumes of 10, 25, 50, 100, 200, 500, and 1000 for tap water and deionized water (dei, 2025). We perform 10 replicate measurements. Our results in Fig. 12a show that as volume increases the magnitude of the capacitance increases, and stabilizes at 500 or higher. We select 500 as the sample volume to use for our main evaluation to increase the separability between liquid classes.
6.1.3. Temporal noise characterization and reduction
Here, we characterize the touchscreen’s sensor noise over time. We note that because the touchscreen is primarily designed to detect touch locations rather than precise capacitance values, it is inherently not optimized for low-noise measurements. To characterize the noise profile, we record the tablet capacitance readings over the course of 30 seconds without anything on the screen. We show in Fig. 13a the standard deviation per pixel across the full screen. We can observe the standard deviation is high in the center of the screen, and lower at the sides, the average standard deviation is 8. To reduce the effect of noise we perform averaging across each frame. Fig. 13b shows the effect of averaging across 20 frames (12 seconds), with an average standard deviation across the screen of 2.8. The relationship between averaging time and the average standard deviation across the screen is shown in Fig. 13c, the plot shows the theoretical relationship of , where is the number of frames being averaged over as well as the measured decrease in standard deviation over 30 seconds. The plot shows that the measured readings decrease with a trend closely following the theoretical model.
6.1.4. Effect of Interfacial Layers
Here we consider the effect of two types of interfacial layers between the touchscreen glass and the liquid: thin-films and containers.
Thin-film Insulating Layer. We examine the effect of adding a thin-film layer on top of the tablet touchscreen on the measured capacitance (Fig. 12b). In principle having the least amount of separation between the liquid sample between the sample and the electrodes will yield the highest capacitance changes. However, one may wish to add an insulating layer when testing viscous liquids such as oil or milk, or biomedical liquid samples such as DNA.
We used three films: food wrap () (foo, 2025), Kapton tape () (kap, 2025), and PET plastic () (pet, 2025) (similar to material used in overhead transparency films (ove, 2025)), where is the thickness of the film. We show the centroid capacitance for different volume levels of water. The results in Fig. 14a show the capacitance change for thinner films is larger than with thicker films, with differences showing starting at 50 and becoming more pronounced at larger volumes up to 1000 .
The reason for this is that when an additional dielectric layer is added the total capacitance can be modeled as a series of capacitors representing the glass screen, the thin film, and the liquid sample:
Because capacitances in series combine reciprocally, the presence of the thin film reduces the overall effective capacitance and attenuates the measured signal.
Through-Container Sensing. Given that the electric field is able to pass through thin films with up to as in the case of PET plastic, we next examine the effects of through-container sensing of liquids in containers (Fig. 14). To do this, we select three different containers: a plastic drinking cup (), a beaker (), and a vial (), and filled them with water until they reached the 2 cm mark and placed them on the touchscreen. The measured heatmaps reveal four points:
The largest capacitance changes occur along the container’s rim. This occurs because the base is slightly concave, causing only the perimeter to make firm contact with the screen while the center remains slightly elevated.
The perimeter exhibits negative capacitance (positive in device units), resembling the response from a finger touch, indicating that conductive coupling dominates here rather than permittivity/dielectric effects as seen with thin films.
In contrast, at the center of the container, capacitance increases (negative in device units), which is similar to the response observed with thin films dominated by the permittivity effect.
There are fringing effects at the four corners around the cup even at electrodes where the cup makes no contact with the screen. This occurs because the electric field lines are not perfectly confined between adjacent electrodes and extend outward into the surrounding region. When a large container perturbs the field, these fringing lines couple capacitively to the container and the liquid inside it.
6.2. Feature extraction and learning
Based on the findings from our empirical analysis and modeling, we design a feature extraction and learning-based pipeline for downstream liquid detection and classification applications.
6.2.1. Automatic droplet identification
To automatically identify the location of a drop, we use a -score-based detection algorithm. For each frame, the algorithm computes the mean and standard deviation of all pixel values and applies a negative threshold defined as , where is the mean, is the standard deviation, and is a user-configurable parameter set to 2, reflecting the decreased capacitance effect (in device units) of a droplet on the screen. Cells that fall below this threshold are considered as candidate droplet locations.
Our algorithm then searches for the connected component (cells) around the candidate locations. We consider cells as connected if they share an edge of a corner, including diagonal neighbors (8-connectivity). This search produces a set of candidate regions for the droplet.
To filter out noise, the algorithm removes regions smaller than a minimum size threshold. When considering droplets that are 50 or smaller which only occupy one cell, we set this threshold to 1. The user can set in the app the volume of the drop they wish to deposit, and using prior information about how large of a surface area the droplet could probably spread, the app can set a conservative minimum size threshold.
Finally, regions are further filtered based on geometric constraints. Specifically, we assume that droplets are approximately elliptical in shape and that the ratio between their major and minor axes do not differ significantly. We discard candidate regions if the difference between length and width does exceeds 2 pixels. This approach effectively isolates compact, roughly elliptical regions corresponding to droplet locations.
Instead of running this algorithm continuously, we only do so when a droplet has been detected to have been deposited on the screen. We do this by maintaining a log of frame to frame changes using the metric: , where is the heatmap at timestep . When exceeds a running threshold , we mark that frame as being when a drop has been deposited. We set the sensitivity parameter to be 2.
6.2.2. Capturing liquid spreading behavior
Prior works (Horstmann et al., 2021) on liquid sensing using capacitive sensing have focused on reporting the capacitance value at the centroid of the liquid deposit location. We note that using this alone is typically insufficient to obtain high classification accuracies at distinguishing between different liquids and classifications, and error bars between different classes overlap. This is due at least in part because a single centroid value does not capture spatial variations in how a liquid spreads.
Our approach instead captures capacitance information in the patch around the centroid which covers a sufficient neighborhood to capture spatial variations caused by different liquid drops while keeping input size computationally manageable. This spatial information implicitly captures the effects of surface tension which affect how the liquid spreads across the surfaces and interacts with neighboring electrodes. For example as shown in Fig. 7, when comparing water () and isopropyl alcohol (), alcohol spreads and flattens more easily. So for the same volume, alcohol occupies a larger surface area and changes the capacitance in a larger neighborhood compared to water.
We designed a compact convolutional neural network to comprehensively extract all these features effectively, specifically designed for single-frame classification. The network includes two convolutional layers with 3×3 kernels, increasing channel depth from 1 to 32 and then 64, each followed by max pooling. Dropout is applied after the convolutional blocks to prevent overfitting. The flattened features pass through a fully connected layer with ReLU activation, another dropout layer, and an output layer for classification. This lightweight model effectively extracts spatial features from individual frames, enabling reliable detection of adulteration types and other liquid categories.
7. Implementation
All of our experiments are performed on the Samsung Galaxy Note 10.1 (2014, SM-P600) (not, 2025), which is rooted to provide super-user capabilities to access the underlying mutual capacitance values. The tablet uses an Atmel maXTouch mxt1664S display controller, which can be accessed using Atmel’s open-source command-line utility, mxt-app (atm, 2025). The Atmel utility reads the mutual capacitance measurement data via I2C communication, polling the register in a loop. It can read the matrices of the capacitance references and deltas which is the difference between the measured and reference capacitance values. The fastest it is able to read is at a period of 0.6 s between frames. The software produces a capacitance matrix of size 52 32 cells, given that the tablet screen has a physical size of 21.6 13.5 cm, this indicates an electrode pitch of 4.2 mm. It produces deltas and references of the heatmaps.
For preprocessing, we measure the reference capacitance heatmap once at the start of each measurement session. We then continuously measure the delta values at each frame. The measured capacitance at frame is obtained as:
As the dielectric system of the touchscreen also includes the protective glass screen on top and air gaps (which all have relative permittivity ), which also affects the electric field, we remove this baseline by subtracting the first measured heatmap from all subsequent heatmaps, giving:
7.1. Software infrastructure
We developed three pieces of software to assist with the design of our system:
Firmware Patch for Capacitance Access. The current version of the utility provided by Atmel did not work on our version of the tablet as it had strict hardware validation checks for optional controller objects that would cause the app to terminate. We wrote a software patch to swallow these errors, log warnings and use default values instead.
Visualization and Annotation Tool. This is a Python app (Fig. 15) that allows viewing and segmenting the droplets from the heatmap file for downstream use in our machine learning algorithms. It provides features for automatic and manual labeling of droplet location, real-time adjustment of parameters, animation of the heatmap over time, and printing of different statistics about the selected droplet regions.
Real-Time Tablet Display App. This is an Android app that calls the mxt-app utility to write the capacitance values to disk, reads the recordings from disk and displays the values for each cell to the tablet screen. It implements computer vision thresholding algorithms to identify candidate droplet deposition locations, marks them, and runs them through a designated classifier to output in real-time the classification result on the screen. We find that a conservatively safe polling interval for a full cycle of write and read commands is 2.5 s.
We will open-source this software infrastructure along with an end-to-end tutorial upon publication to enable others to build on top of these utilities.
7.2. Training methodology
We construct a frame-wise training framework in which each time frame from each region is treated as an individual data sample. Compared with region-wise training, which averages all frames within a region and thus loses valuable temporal variation, and sequence-wise training using CNNs combined with LSTM or Transformer architectures, the frame-wise approach makes fuller use of the available data and maximizes the effective dataset size. Moreover, since our measurements contain limited temporal dynamics, employing LSTM or Transformer models to extract temporal features would not provide significant benefits.
Specifically, we extract 50 time frames from each region, thereby expanding the dataset size by a factor of 50. The classifier is trained and evaluated using 5-fold cross-validation to ensure robust performance, with training parameters set to 50 epochs, batch size 16, learning rate 0.001, and a ReduceLROnPlateau scheduler to mitigate overfitting.
8. Evaluation
8.1. Adulteration detection
Here we evaluate if our system is able to distinguish adulterated liquids and we outline the three scenarios below:
Soda spiking. Here we consider the scenario of a soft drink being spiked. To do this, we use as our base soda Coke, and spike it with different concentrations of industrial ethanol, a form of alcohol unsuitable for human consumption. Ethanol is added at concentrations of 10, 20, 40, 60, and 80% by volume. This scenario simulates intentional contamination with toxic substances. The dataset includes 77 unadulterated and 377 adulterated samples.
Wine adulteration. Next, we investigate the adulteration of wine, which already contains alcohol. Here, the base liquid is red wine, to which industrial ethanol is added in increments of 20%. Compared to soda spiking, this scenario is more challenging because the background alcohol content in wine reduces the relative change caused by adulteration. In total, 88 unadulterated and 231 adulterated samples are collected.
Milk adulteration. We further explore milk adulteration, a common issue in dairy products where substances like salt and detergent are sometimes added to mimic the taste or appearance of pure milk. Salt can mask dilution, while detergent creates artificial froth but poses health risks. To simulate such cases, we mix 2 g of salt into 20 mL of milk and 2 mL of detergent into 20 mL of milk. The total numbers of unadulterated and adulterated samples are 44 and 74 respectively.
Our results in Fig. 16a show the precision and recall at detecting the adulterated class. We note that these metrics of interest better capture the expected performance compared to accuracy because they emphasize the model’s ability to correctly identify adulterated samples even if it is at the cost of occasional false positives. In our application, the occasional false positive is acceptable as it may simply warrant re-testing, or in the worse case, discarding of the liquid sample. The converse is not acceptable as missing an adulterated product poses a greater risk than a false alarm. Our results show our precision and recall is above 90% for the three adulteration detection tasks.
We show the confusion matrix performance for the three tasks in Fig. 16b–d, the results show that the overall accuracy of these classifiers are all above 96%. In the case of soda spiking the rate of false positives is less than 1%. For wine adulteration detection, we note this is a more challenging case as wine already contains alcohol so when it is contaminated with certain concentrations of alcohol, it is more challenging to detect the adulterated sample, however the overall accuracy is 96.1%. In the case of milk adulteration, the overall classification accuracy reaches 98.1%. Notably, for detergent adulteration, which poses genuine health risks to humans, the false detection rate is 0.0%, indicating that the system can reliably distinguish even subtle forms of contamination. These results collectively demonstrate that our method provides a robust and reliable approach for detecting adulteration in various beverages and liquid foods.
8.2. Concentration detection
Here, we evaluate the extent to which our system could be used to distinguish between different concentrations of liquids. We consider distinguishing between different concentrations of DNA and NaCl salt water using the same concentration breakdown as in Fig. 8. We also perform this for ethanol with slightly different concentrations, though still covering the full 0 to 100% range. The goal of this particular analysis is to determine the granularity at which these concentrations can be detected.
We find that in the case of DNA and salt water (Fig. 17a,c) our system is primarily effective as a binary detector of trace concentrations of liquid. Specifically in the case of DNA it is able to consider concentration of 0–20 ng/ as a low concentration and 40 ng/ as a high concentration and distinguish between those, essentially functioning as a detector with a threshold of 20 ng/, yielding an overall detection accuracy of 95.6%. Similarly, in the case of salt water, the touchscreen is able to function a detector of the liquid with a threshold of M with an overall detection accuracy of 95.7%. In the case of ethanol, we find that our system is able to distinguish between six different concentrations of alcohol specifically 0, 10, 30, 50, 80, 100% with an overall accuracy of 93% (Fig. 17b).
8.3. Through-container sensing
We further evaluate the system’s performance in through-container liquid concentration and adulteration detection, as directly dispensing liquids onto the touchscreen may not be practical in real-world use. In such cases, the base of the container introduces an additional dielectric layer that partially attenuates the capacitive response from the liquid in it.
As shown in Fig. 14, placing a container on the screen produces a complex spatial response pattern, characterized by a strong positive capacitive change along the container’s rim, negative values at the center, and fringing effects near the corners. We observe that the rim region exhibits the most pronounced and consistent change across different liquids and container types. Therefore, we use the mean positive value along the rim as the primary feature for distinguishing between containers. We also used a 9 oz plastic cup as we find it generates the strongest capacitive response on the touchscreen due to its thin base.
We evaluated the through-container sensing capability to distinguish between different types of liquids. Three classification tasks were conducted:
Basic liquid classification: Tap water, deionized water, and pure ethanol were tested to verify that the model can distinguish liquids with large differences in capacitance. The number of samples for tap water, deionized water, and pure ethanol were 18, 20, and 16, respectively.
Soft drink adulteration detection: To evaluate the model’s ability to identify beverage adulteration, we used soda as the base liquid and adulterated it with 20% industrial ethanol. Both unadulterated and adulterated soda classes contained 16 samples.
NaCl concentration detection: To test sensitivity to liquid concentration, we compared deionized water, 0.001% NaCl solution, and 0.01% NaCl solution, each with 20 samples.
Initially, we applied the same CNN model used in the drop-level liquid classification. However, due to the larger input dimension (approximately 13×13 pixels compared to 6×6 or 8×8 previously) and the more complex spatial pattern introduced by the container, the CNN struggled to learn discriminative features effectively, likely also affected by the limited dataset size. To address this, we adopted a feature-based random forest classifier using manually selected statistics from the region’s positive responses: the mean, median, and 75th percentile of positive values. We show the positive mean of different liquids tested in our through-container evaluation in Fig. 18.
As shown in Fig. 19, through-container sensing achieved 95.8% accuracy in the tap water–DI water–ethanol task, 90.9% in soda adulteration detection, and 85.7% in NaCl concentration classification. We note that these accuracies are slightly lower than when the liquid drop is directly on the screen, however this is expected as signal attenuation from the container’s base can create more variation. However we note that the performance is still high exceeding 90% on two of the tasks.
8.4. Benchmarks
Measurement stability over time. We evaluate the stability of capacitance readings for the priming drop and the test liquid samples on the tablet screen. We show in Fig. 20a that the measured capacitance at the centroid of the priming drop remains stable across the measured period of ten minutes with a mean and standard deviation of 376 7 device units. We also perform this measurement for water samples of different volumes across an 10 minute period and show in Fig. 20b that these values remain stable across the measured period with standard deviations ranging from 6 to 10 device units across the measurement period. There is no substantial relationship between the standard deviation of the measurements with the volume of the liquid drops. We note that stability across these time scales is sufficient for averaging the effects of noise and for recording a measurement.
Effect of temperature. We examine the effect of temperature on measured capacitance. We perform this experiment using tap water and deionized water. We heated up the liquid samples in a tabletop water bath (Fig. 21a) (ivy, 2025) from 30–90∘C. The results in Fig. 21b show that the measured capacitance increases for both liquids. We note that the reason is different for both liquids tested. In the case of tap water which has more ions than deionized water, the ions mobility will increase which will increase the conductivity, and thus a larger capacitance. In contrast, with deionized water, where there are no ions, the capacitance increase is due to changes in the permittivity which increases the capacitance. We note that in this experiment, the measured capacitance for the tap water increases to approximately -800 device units, which is the empirical limit of the system.
Effect of container type and liquid volume. We compare three types of containers: 9 oz plastic cups (pla, 2025), 25 mL glass vials, and 5 mL glass vials (via, 2025), chosen for their relatively flat and thin bottoms, which allow stronger capacitive coupling with the touchscreen. To investigate the effect of liquid volume, we test tap water at heights of 12, 24, and 36 mm in the 9 oz plastic cup and 25 mL vial. For the smaller 5 mL vial, only the 12 mm water level is feasible. Additionally, we test a small 5 mL water volume in the plastic cup to isolate the effect of liquid quantity from container type.
The results in Fig. 22 show that, for a given container, changing the water level or liquid amount does not significantly alter the positive mean response. This suggests that the dominant factor influencing the touchscreen’s capacitive response is likely the container’s base thickness and material, rather than the volume of liquid inside. Since the 9 oz plastic cup generated the strongest capacitive response on the touchscreen, we selected it as the standard container for subsequent experiments. We used 23 mL of liquid for testing, although 5 mL produced a comparable response, 23 mL represents a more practical amount for everyday use.
Effect of device tilt. As the user may not always have their tablet flat on a surface, particularly in field conditions outdoors, the effect of uneven surfaces on our system. To do this, we placed the tablet on a motorized tilt platform (til, 2025) (Fig. 23a) and adhered the tablet to it using some putty. We deposited ten drops of water and isopropyl alcohol (IPA) onto the surface and varied the tilt angle at increments of approximately 1 to 3∘. We show in Fig. 23b how the droplets displace from their original position as tilt angle increases. In the case of water, which has a surface tension of , the droplets start moving at 5∘ and this displacement increases till its roll-off angle of 11.5∘ when it slides off the tablet. In contrast, IPA, which has a lower surface tension of , has a much higher rate of displacement with a roll-off angle of 2∘. While water had a displacement of approximately 4 cells at the roll-off angle, alcohol had a displacement around three times higher at 12 cells. The behavior of the drops at the rest and roll-off angles are visualized in Fig. 23c.
Ablation study. We perform an ablation study on the liquid adulteration datasets to examine the effect of different system design choices on performance (Fig. 24). Specifically, we examine the effect of number of training samples, region size around the centroid, and machine learning model on system performance. For these experiments, we consider a subset of the adulteration classes: for soda spiking and wine adulteration we compared to ethanol of 10, 20, 30% as a matter of computational efficiency to examine the overall effect of the system parameters. Here we used 500 drop sizes.
For the training dataset size (Fig. 24a) we evaluated the affect of changing the size of the training set from 0, 25, 50, 75, 100% of the original size. We find that as expected this increases the accuracies with diminishing returns. In the soda and wine dataset, we find that there are no substantial performance increases beyond approximately 200 training samples.
For the region size (Fig. 24b), we find that accuracy increases steadily as the region size increases from a patch to a patch. In the case we use a fully-connected network instead of a CNN. For larger classes we use an adaptive CNN where for and classes there are only two convolutional layers without any pooling, while for larger patch sizes of to there is a single max pooling layer, and for there are 2 pooling layers. This result is expected as the screen area covered by a drop can expand to affect the electric field of a grid, and a larger grid captures this information.
Finally, we examined the effect of model architecture on system performance (Fig. 24c). We find that using a CNN only, LSTM only, CNN+LSTM and CNN+Transformer provide performance within a similar range of 91 to 94%. Tree-based models such as XGBoost and Random Forest provide lowered accuracies of 79 to 81%. The results suggest that global features about the region, can be captured by the tree-based models. However, deep learning modules better capture the non-linear relationship (Sec. 5) between classes, with the CNN and LSTM producing higher accuracies. As CNN alone and LSTM alone are able to produce high accuracies, it suggests that non-linear feature combinations are sufficient to perform classification.
8.5. User study
We conducted an user study determined exempt by our IRB (STUDY2025_00000371) and recruited 9 participants by word of mouth from our institution. All subjects were PhDs students in computer science and engineering. We note that prior literature (Nielsen and Landauer, 1993) indicates testing on 5–8 participants is sufficient to uncover 85% of usability issues in a system.
Study design. The goal of this user study was to evaluate our system’s ability to detect drops deposited by lay users, as well as the accuracy and time for user’s to deposit drops on the screen. We also assess the system’s workload and user’s subjective evaluation of the system. At the start of the study, we showed participants a demo of the real-time app illustrating that placing drops of liquids would change the capacitance heatmap. We then gave them a tutorial for how to use a pipette for the purposes of the usability study as none of them had used a pipette before. However, we note that lay users could always use eye drops which draw up a constant amount of fluid each time, and is effective at depositing the priming drop. We showed them the app interface (Fig. 25a,b) which displayed an area to deposit the priming drop, and grid lines indicating individual electrodes, the center of which users had deposit the liquid sample of interest. The software marked a set of 24 cells split into 3 rows and 8 columns where the users had to deposit a drop. For the first row, users needed to deposit 8 drops of 10 across the 8 columns. This is repeated for the second and third rows where they had to deposit 25 and 50 samples.
Results. Below we describe the results of our user study:
Drop detection rate. We computed how many drops were detected by our software (Sec. 6.2.1). In the 10 case, we find that for the 8 drops 9 users = 72 drops, 67 of 72 drops (93%) were detected by our algorithm, while for the 25 and 50 case, 71 of 72 (99%) drops were detected by our algorithm. We note that our system provides visual feedback by highlighting the detected drop area, allowing it to prompt the user to deposit the drop within a specified time window and re-prompt if no drop is detected.
Accuracy of deposition location. We computed the centroid of each drop and measured its deviation from the marked target cell. We find that across all but one detected drops, they were deposited precisely at the marked location. The single drop not correctly deposited only deviated from the correct location by one cell.
Time to deposit drops. We measured the duration to deposit a row of 8 drops of 10, 25, and 50 . We plot the average time to deposit a single drop in Fig. 25c and show that for the 10 case users took on average 8.6 2.0 s to deposit a drop which was approximately 1 s longer than the time to deposit 25 and 50 drops which took 7.5 2.4 s and 7.6 1.9 s on average respectively. This is likely because users found it more difficult to control the smaller 10 drops which required finer hand movements to release accurately onto the marked positions on the grid. We also show the time for a trained user to deposit the drop which ranged from 4.5 to 5.5 s, which is around 2–3 s faster than the average untrained user. Overall, we note that the time to deposit liquids onto the screen was relatively quick.
Post-task questionnaire. Upon completing the task, we evaluated the subjective workload of using our system through the NASA Task Load Index (TLX) (Hart, 2006) where users were asked to assess workload across five scales which each had 21 gradations, they were asked:
How mentally demanding was the task?
How physically demanding was the task?
How hurried or rushed was the pace of the task?
How hard did you have to work to accomplish your level of performance?
How insecure, discouraged, irritated, stressed, and annoyed were you?
The 21-point scale was categorized into low (0–6), medium (7–13), or high (14–21) effort categories. The results (Fig. 25d) show that across all six scales, the average rating across the five task dimensions ranged from 3 to 6 which were within the low workload category. We note that some users rated some workload demands as medium and hard. We suspect this more likely reflects the use of a pipette to draw up and deposit liquid droplets, but we note that this procedure would be similar to using pipetting samples for use on lab-grade instrument. This scenario also reflects a more challenging use case of our system, as users can also use an eye dropper with a larger amount of liquid such as 500 which can be detected by our system. While our study focused on users placing the sample at a fairly precise location on the screen, our system is capable of sensing liquid through-containers as well.
Users were also asked three further questions, specifically:
How successful were you in accomplishing what you were asked to do?
0 = Perfect, 21 = Failure -
How satisfied are you with the system’s overall ability to perform liquid sensing?
Likert scale of 1 = Very dissatisfied, 5 = Very satisfied -
How much would you be willing to pay for this system?
Users’ subjective evaluation of their success at the task as on average 6 4, suggesting that on average users felt generally successful, but with minor issues (Fig. 25d). The users’ overall satisfaction with the system (Fig. 25e) was on average 4.1 0.7 which leans towards satisfied. Finally, users indicated they were on average willing to pay $21.05 for the system, we note that this number also implicitly reflects spending preferences of our subject population.
9. Discussion
9.1. Limitations
Detectable liquids. As capacitance sensing depends on a material’s conductivity and permittivity, liquids with low values for both, such as oils, are inherently difficult to detect. In particular oils have relative permittivity values that typically range from 1–2, which is substantially lower than water, which is 80 (Dhekne et al., 2018). In preliminary testing, we find that the screen is able to register drops of oil, however, the measured capacitance values are low. This is however a general limitation of capacitance-based sensing. Our system is therefore not intended as a universal liquid sensor. Instead, primarily targeted towards detection of liquids with contrasting dielectric values or concentration detection.
Geometry and thickness of container base. The containers that we selected for our through-container sensing experiments had flat and thin bases to enable capacitive coupling between the touchscreen electrodes and the liquid sample. However, we note that containers with other base geometries—such as curved or irregular bottoms—can produce nonuniform or distorted response patterns on the screen, while thicker bases tend to attenuate the signal and may reduce it below the touchscreen’s detection threshold.
Broader device integration. A limitation of this work is that the system has only been evaluated on a single device. This is primarily because manufacturers typically do not provide public access to raw capacitance heatmap data, which is typically available only through research collaborations. However, our work represents a first step towards what could be made possible through this sensing modality.
We note that the field of ubiquitous computing and mobile systems has a long and rich history of recognizing the value of research leveraging specialized or restricted hardware capabilities available to only a fixed range of devices or requiring some level of software or hardware modifications, including near-infrared cameras (Barry et al., 2022), LiDAR (Chan et al., 2022b), high-frame-rate capacitive sensing displays (Wang et al., 2019), rooted devices with custom kernels to increase IMU sampling rates (Laput et al., 2016), thermal cameras (Adhikary et al., 2024), or in-ear microphones on earbuds (Cao et al., 2024; Butkow et al., 2023; Zhang et al., 2025). In the same spirit, our work illustrates a plausible technical pathway for how tablet touchscreens could be used for liquid sensing.
Experimental control factors. We note that different physical factors affect the screen capacitance and need to be controlled to ensure a fair comparison between liquid samples. First, liquid capacitance varies based on temperature and humidity, Second, sample volume affects the heatmap signature of the drop given that larger volumes spread over a larger area. Thus standardized environmental conditions and operating procedures would be needed for accurate and repeatable results, however this requirements are no different than for lab-grade devices.
9.2. Future directions
Leveraging onboard actuators to measure other liquid properties. Future work could could focus on investigating if other onboard motors on the tablet such as the vibration motor and the backlight LCD/OLED could be used for measuring other liquid properties:
Vibration motor. In initial testing, we observed that when the tablet is balanced on top of a surface such as the motorized tilt platform in Fig. 23, the vibration motor generates sufficient force to agitate and gradually move small objects such as a 10 g weight (wei, 2025), across the screen. This occurs when the material and geometric properties of the object resonates with the motor’s fundamental frequency. We note that the object does not move when the tablet is placed flat on the table, which dampens the effect of the vibration.
We observed that depositing a drop of liquid on the screen and subsequently placing an object on it resulted in the liquid being smeared across the surface as the object moved. Further investigation is required to determine (1) if these smearing signatures are correlated to other liquid properties such as viscosity and surface tension, (2) how to design or select objects which are maximally resonant with the motor (3) the effect of the object’s initial position with respect to the motor on its trajectory. Given that this is an independent idea with its own set of technical challenges, we leave this as a future direction of investigation.
Display backlight. Modern touchscreen displays typically provide backlight intensities of 500–1000 nits (nit, 2025) with up to 2000 nits in some cases. This illumination could potentially serve as an integrated optical excitation source for modulating the physical properties of liquid or solid samples placed on the screen. Such a capability may be particularly useful for characterizing photoconductive materials whose capacitance varies with light exposure.
An alternative formulation of such a system would be to have an external optical excitation source that can sweep over a larger optical wavelength range such as ultraviolet, visible-light, and near-infrared to impact multiple liquid samples on the screen simultaneously while using the screen as a reader. We leave this as a future direction for investigation.
Sensing of soil and solids. Given the sensing capabilities of tablet touchscreens, characterizing the capacitive properties of different solids may be a useful further direction. In particular measuring soil moisture levels could be an interesting direction given that soil analyzers (soi, 2025a, b) often rely on capacitance-based sensing using parallel plate capacitors. We note that touchscreens are also sensitive enough to detect the presence of conductive materials such as copper (conductive tape), coins (nickel, copper, and zinc), and keys (brass, nickel) register, with these capabilities being used in prior works (Mayer et al., 2021; Villar et al., 2018; Xiao et al., 2016; Pedersen and Hornbæk, 2011; Voelker et al., 2013, 2015; Chan et al., 2012; Kratz et al., 2011; Liang et al., 2014; Schmitz et al., 2017) on enabling tangible user interfaces and widgets that can interact with the touchscreen.